Three particles of equal masses m are placed at the vertices of an equilateral triangle of 

side 2.0 m. Determine the centre of mass of this system.


1
Expert's answer
2021-01-19T12:25:24-0500

Answer

Suppose one of corner of triangle is at origin.

equilateral triangle mean internal angle is 60degree.

So all vertex of triangle coordinates can be written as

(0, 0), (1, 1.73), (2, 0)

The centre of mass of this system

"X_{com}=\frac{m(x_1+x_2+x_3) }{3m}"

"X_{com}=1"

Now

"Y_{com}=\frac{m(y_1+y_2+y_3) }{3m}" =0.57